OpenStudy (anonymous):

A waiter earned $23 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 15 bills.

OpenStudy (johnweldon1993):

We have a little system of equations here let x represent the number of 1 dollar bills and 'y' represent the number of 5 dollar bills we know they got 15 bills total so x + y = 15 now we also know they earned 23 dollars total...so using the value of the money TIMES how much they got....we have 1x + 5y = 23 so how can we solve that system of equations? x + y =15 1x + 5y = 23

OpenStudy (texaschic101):

good explaining @johnweldon1993 .....you couldn't have explained it any simpler :)

OpenStudy (johnweldon1993):

Thank you @texaschic101 I appreciate that :)

OpenStudy (texaschic101):

just telling the truth :)